OpenStudy (anonymous):

y=5-x and 2x+3y=12 solve by substitution

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

x=3 y=2

OpenStudy (anonymous):

2x + 3(5-x) = 12 2x + 15 - 3x = 12 -x = -3 x= 3 y = 5 -(3) = 2

OpenStudy (anonymous):

2x+3y=12________________1 y=5-x___________2 substitute 2 in 1 2x+3(5-x)=12 2x+15-3x=12 -x=-3 x=3 and y=5-3 y=2 (x,y)=(3,2)